OpenStudy (anonymous):

How do you solve 2sin^2 (theta) + cos(theta)=2?

OpenStudy (anonymous):

try substituting (sin^x)^2=1-(cos^x)^2

OpenStudy (anonymous):

60 OR 90 DEGREE

OpenStudy (anonymous):

\[2\sin^2\theta+\cos \theta=2\] We know that: \[\sin^2\theta = (1-\cos^2\theta)\] So we can substitute and get: \[2(1-\cos^2\theta)+\cos \theta = 2\]